# Rethinking the Role of Prompting Strategies in LLM Test-Time Scaling: A Perspective of Probability Theory

### v2

@@ 2.4 Formal Expression
$s.t.~{}\sum_{x\in\mathfrak{D}}\mathcal{C}(x\,|\,\mathcal{M};\mathbf{P}_{i};N)\leq O.\vspace{-2pt}$ (3)
@@ 4 Why CoT Performs worse with Lower $N$ while better with Larger $N$?
Figure 3: Illustration of the two reasons why CoT sometimes performs worse with lower $N$ while better with larger $N$. Left: CoT has more easy questions and fewer hard questions. For example, the probability distribution of L2M is $\{0.4,0.5,0.1,0.0,0.0\}$ (hard question), and $\{0.3,0.2,0.2,0.2,0.1\}$ (easy question) for CoT. Although L2M has higher pass@1 accuracy, its accuracy reduces until 0% as scaling while CoT increases until 100%. Right: CoT is less likely to be affected by wrong answers due to their relatively uniform distribution. The probability of obtaining the right answer ${a}_{1}$ grows more rapidly as increasing $N$.
@@ 4.2 CoT is Less Likely to be Affected by Wrong Answers
The second reason for this phenomenon is that CoT is less likely to be affected by wrong answers. $\text{Pr}({a}_{1}|\mathbf{P}_{i};N)$ is not a function of only the probability $p_{i,1}$ of the right answer ${a}_{1}$, but also related to the probability distribution of other wrong answers. According to Theorem 4, even if $p_{i,1}>p_{i^{\prime},1}$, i.e., $\text{Pr}({a}_{1}|\mathbf{P}_{i};N=1)>\text{Pr}({a}_{1}|\mathbf{P}_{i^{\prime}};N=1)$, there still may exist an $N_{0}$ that $\text{Pr}({a}_{1}|\mathbf{P}_{i};N=N_{0})>\text{Pr}({a}_{1}|\mathbf{P}_{i^{\prime}};N=N_{0})$. Considering a question $x$ in GSM8K as an example and ${a}_{1}$ is the correct answer, the result probability distribution of $\mathbf{P}_{i}=\text{SBP}$ is $\{0.64,0.35,0.01\}$, and $\{0.6,0.2,0.2\}$ for $\mathbf{P}_{i^{\prime}}=\text{CoT}$, which satisfies the condition in Theorem 4. According to Lemma 1, $\text{Pr}({a}_{1}|\mathbf{P}_{i};N=1)=0.640>\text{Pr}({a}_{1}|\mathbf{P}_{i^{\prime}};N=1)=0.600$, $\text{Pr}({a}_{1}|\mathbf{P}_{i};N=3)=0.709>\text{Pr}({a}_{1}|\mathbf{P}_{i^{\prime}};N=3)=0.696$, while $\text{Pr}({a}_{1}|\mathbf{P}_{i};N=5)=0.757<\text{Pr}({a}_{1}|\mathbf{P}_{i^{\prime}};N=5)=0.769$. This means, although complicated prompting strategies may have higher pass@1 accuracy, they are easier to be affected by wrong answers. In contrast, simple CoT has a relatively flat distribution on wrong answers, thus making it focus more on the correct answer, which makes it more rapidly improve performance in easy questions and more slowly reduce accuracy in hard questions as increasing $N$, as shown in Figure 3. We record the quantity of such questions for each two prompting strategies and display the results of Qwen2.5-7B-Instruct and LLaMA-3-8B-Instruct in Table 2. If CoT is $\mathbf{P}_{i^{\prime}}$, there are the most data satisfying Theorem 4. If CoT is $\mathbf{P}_{i}$, there are the least such questions. These demonstrate that CoT has greater potential to significantly increase scaling performance compared with other strategies.
@@ Appendix B Proofs
$\begin{matrix}\!\!\!\!\!\!\!\!\text{Pr}\,({a}_{1}|\mathbf{P}_{i})=\text{Pr}\,(\mathbf{x}_{1}>M)+\sum_{|\mathcal{J}|=1}^{m-1}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\vspace{10pt}\\ \dfrac{\text{Pr}\,(\mathbf{x}_{1}=\mathbf{x}_{j}>\mathbf{x}_{q}~{}{for~{}all}~{}j\in\mathcal{J},q\notin\{1\}\cup\mathcal{J})}{|\mathcal{J}|+1},\\ \end{matrix}$ (14)
@@
$\vspace{10pt}\begin{split}&\frac{\left(lk+1+\sum_{j=l+1}^{m}k_{j}\right)!}{(k+1)\cdot(k!)^{l}\cdot\prod_{j=l+1}^{m}{(k_{j}!)}}p_{1}^{k+1}p_{2}^{k}\cdots p_{l}^{k}p_{l+1}^{k_{l+1}}p_{l+2}^{k_{l+2}}\cdots p_{m}^{k_{m}}\\ -&p_{1}\cdot\frac{1}{l}\cdot\frac{\left(lk+\sum_{j=l+1}^{m}{k_{j}}\right)!}{(k!)^{l}\cdot\prod_{j=l+1}^{m}(k_{j}!)}p_{1}^{k}p_{2}^{k}\cdots p_{l}^{k}p_{l+1}^{k_{l+1}}p_{l+2}^{k_{l+2}}\cdots p_{m}^{k_{m}}\\ -&\sum_{s=2}^{l}{p_{s}\cdot\frac{\left(lk+\sum_{j=l+1}^{m}{k_{j}}\right)!}{\frac{k+1}{k}\cdot(k!)^{l}\cdot\prod_{j=l+1}^{m}(k_{j}!)}}p_{1}^{k+1}p_{2}^{k}\cdots p_{s-1}^{k}p_{s}^{k-1}p_{s+1}^{k}\cdots p_{l}^{k}p_{l+1}^{k_{l+1}}p_{l+2}^{k_{l+2}}\cdots p_{m}^{k_{m}}\\ -&\sum_{t=l+1}^{m}{p_{t}\cdot\frac{\left(lk+\sum_{j=l+1}^{m}{k_{j}}\right)!}{\frac{k+1}{k_{t}}\cdot(k!)^{l}\cdot\prod_{j=l+1}^{m}(k_{j}!)}p_{1}^{k+1}p_{2}^{k}\cdots p_{l}^{k}p_{l+1}^{k_{l+1}}p_{l+2}^{k_{l+2}}\cdots p_{t-1}^{k_{t-1}}p_{t}^{k_{t}-1}p_{t+1}^{k_{t+1}}\cdots p_{m}^{k_{m}}}\\ -&\sum_{s=2}^{l}{p_{s}\cdot\frac{1}{l}\cdot\frac{\left(lk+\sum_{j=l+1}^{m}{k_{j}}\right)!}{(k!)^{l}\cdot\prod_{j=l+1}^{m}(k_{j}!)}p_{1}^{k}p_{2}^{k}\cdots p_{s-1}^{k}p_{s}^{k}p_{s+1}^{k}\cdots p_{l}^{k}p_{l+1}^{k_{l+1}}p_{l+2}^{k_{l+2}}\cdots p_{m}^{k_{m}}}\\ =&\sum_{s=2}^{l}\left[\frac{1}{l}\cdot\frac{\left(lk+\sum_{j=l+1}^{m}{k_{j}}\right)!}{(k!)^{l}\cdot\prod_{j=l+1}^{m}(k_{j}!)}p_{1}^{k}p_{2}^{k}\cdots p_{s-1}^{k}p_{s}^{k}p_{s+1}^{k}\cdots p_{l}^{k}p_{l+1}^{k_{l+1}}p_{l+2}^{k_{l+2}}\cdots p_{m}^{k_{m}}\right]\cdot{(p_{1}-p_{s})}\geq 0\end{split}$ (23)

### v3

@@ 2.4 Formal Expression
$s.t.~\sum_{x\in\mathfrak{D}}\mathcal{C}(x\,|\,\mathcal{M};\mathbf{P}_{i};N)\leq O.\vskip-2.0pt$ (3)
@@ 4 Why CoT Performs worse with Lower $N$ while better with Larger $N$?
[…]
@@ 4.2 CoT is Less Likely to be Affected by Wrong Answers
[…]
@@ Appendix B Proofs
$\begin{matrix}\!\!\!\!\!\!\!\!\text{Pr}\,({a}_{1}|\mathbf{P}_{i})=\text{Pr}\,(\mathbf{x}_{1}>M)+\sum_{|\mathcal{J}|=1}^{m-1}~~~~~~~~~~~~\vskip 10.0pt\\ \dfrac{\text{Pr}\,(\mathbf{x}_{1}=\mathbf{x}_{j}>\mathbf{x}_{q}~{for~all}~j\in\mathcal{J},q\notin\{1\}\cup\mathcal{J})}{|\mathcal{J}|+1},\\ \end{matrix}$ (14)
@@
$\vskip 10.0pt\begin{split}&\frac{\left(lk+1+\sum_{j=l+1}^{m}k_{j}\right)!}{(k+1)\cdot(k!)^{l}\cdot\prod_{j=l+1}^{m}{(k_{j}!)}}p_{1}^{k+1}p_{2}^{k}\cdots p_{l}^{k}p_{l+1}^{k_{l+1}}p_{l+2}^{k_{l+2}}\cdots p_{m}^{k_{m}}\\ -&p_{1}\cdot\frac{1}{l}\cdot\frac{\left(lk+\sum_{j=l+1}^{m}{k_{j}}\right)!}{(k!)^{l}\cdot\prod_{j=l+1}^{m}(k_{j}!)}p_{1}^{k}p_{2}^{k}\cdots p_{l}^{k}p_{l+1}^{k_{l+1}}p_{l+2}^{k_{l+2}}\cdots p_{m}^{k_{m}}\\ -&\sum_{s=2}^{l}{p_{s}\cdot\frac{\left(lk+\sum_{j=l+1}^{m}{k_{j}}\right)!}{\frac{k+1}{k}\cdot(k!)^{l}\cdot\prod_{j=l+1}^{m}(k_{j}!)}}p_{1}^{k+1}p_{2}^{k}\cdots p_{s-1}^{k}p_{s}^{k-1}p_{s+1}^{k}\cdots p_{l}^{k}p_{l+1}^{k_{l+1}}p_{l+2}^{k_{l+2}}\cdots p_{m}^{k_{m}}\\ -&\sum_{t=l+1}^{m}{p_{t}\cdot\frac{\left(lk+\sum_{j=l+1}^{m}{k_{j}}\right)!}{\frac{k+1}{k_{t}}\cdot(k!)^{l}\cdot\prod_{j=l+1}^{m}(k_{j}!)}p_{1}^{k+1}p_{2}^{k}\cdots p_{l}^{k}p_{l+1}^{k_{l+1}}p_{l+2}^{k_{l+2}}\cdots p_{t-1}^{k_{t-1}}p_{t}^{k_{t}-1}p_{t+1}^{k_{t+1}}\cdots p_{m}^{k_{m}}}\\ -&\sum_{s=2}^{l}{p_{s}\cdot\frac{1}{l}\cdot\frac{\left(lk+\sum_{j=l+1}^{m}{k_{j}}\right)!}{(k!)^{l}\cdot\prod_{j=l+1}^{m}(k_{j}!)}p_{1}^{k}p_{2}^{k}\cdots p_{s-1}^{k}p_{s}^{k}p_{s+1}^{k}\cdots p_{l}^{k}p_{l+1}^{k_{l+1}}p_{l+2}^{k_{l+2}}\cdots p_{m}^{k_{m}}}\\ =&\sum_{s=2}^{l}\left[\frac{1}{l}\cdot\frac{\left(lk+\sum_{j=l+1}^{m}{k_{j}}\right)!}{(k!)^{l}\cdot\prod_{j=l+1}^{m}(k_{j}!)}p_{1}^{k}p_{2}^{k}\cdots p_{s-1}^{k}p_{s}^{k}p_{s+1}^{k}\cdots p_{l}^{k}p_{l+1}^{k_{l+1}}p_{l+2}^{k_{l+2}}\cdots p_{m}^{k_{m}}\right]\cdot{(p_{1}-p_{s})}\geq 0\end{split}$ (23)